# Fast Transients from Magnetic Disks Around Non-Spinning Collapsar Black Holes

### v1

@@ 4.1 Explosion physics
The bottom panel of Fig. 6 displays the binned mass (red) and energy (blue) distributions of the gas as a function of the dimensionless velocity $\beta=v/c$. The unbound mass satisfies $M_{\rm ub}\sim 10^{-10\beta}$, whereas the energy remains roughly uniform at $\beta\lesssim 0.3$ after which it plummets. Such a uniform energy distribution is similar to those associated with the slower cocoon regions formed by the interaction of relativistic jets with the star (Gottlieb et al., 2021; Eisenberg et al., 2022). This suggests that the observational signatures, which we discuss next, may be similar between disk-powered and jet–cocoon-powered sub-relativistic outflows.
@@ 5 Conclusions
The disk outflows are driven by a comparable contribution of vertical magnetic field lines threading the accretion disk (BP jets) and magnetic flux eruptions from regions of high magnetic pressure near the BH. Feedback from the cocoon, which is generated by the disk winds, induces a disk tilt that causes the outflows to wobble. With an isotropic-equivalent energy of $E_{\rm iso}\approx 10^{52}\,\,{\rm erg}$, these outflows explode the star, leaving merely $\sim 1\,\,{\rm M_{\odot}}$ for BH mass growth. After breakout from the star, the disk outflows will generate multiple types of emission across the spectrum: shock breakout, cooling envelope, ${}^{56}{\rm Ni}$ decay and synchrotron emission. We find that their cooling emission will be detectable out to hundreds of Mpc by ULTRASAT and $D>{\rm Gpc}$ by UVEX and the Rubin Observatory. The detection rate of these events will provide valuable constraints on accretion disk physics and its formation rate in CCSNe.
The UV/optical cooling emission will be followed by ${}^{56}{\rm Ni}$ decay emission. Hydrodynamic simulations by Fujibayashi et al. (2023, 2024) suggest that the ${}^{56}{\rm Ni}$ emission is expected to last a few weeks, however this result might change for magnetic disks [^2] Our hydrodynamic disks fail to launch such powerful outflows (see §3). This discrepancy may arise from the enforced axisymmetry in Fujibayashi et al. (2023, 2024), which requires verification through 3D calculations.. Additionally, recent neutrino–GRMHD simulations of collapsars indicate that strong magnetic fields in the accretion disk can facilitate the ejection of heavy $r$-process elements from the disk (Issa et al., 2024). This could potentially trigger a rapid “kilonova” emission and alter the ${}^{56}{\rm Ni}$ yields compared to those found in hydrodynamic simulations. Future work incorporating neutrino transport in GRMHD simulations is necessary to investigate the impact of disk cooling on our results, and model all three optical signals (cooling, kilonova and 56Ni decay) from first principles.

### v2

@@ 4.1 Explosion physics
The bottom panel of Fig. 6 displays the binned mass (red) and total energy (blue) distributions of the gas as a function of the dimensionless velocity $\beta=v/c$. The unbound mass satisfies $M_{\rm ub}\sim 10^{-10\beta}$, with energy displaying a similar monotonic decline with $\beta$. This decaying energy distribution contrasts with those associated with the non-relativistic portion of the cocoon formed by the interaction between relativistic jets and the star (Gottlieb et al., 2021; Eisenberg et al., 2022). This difference suggests that the observational signatures, discussed in the following section, likely vary between disk-powered and jet-cocoon-powered sub-relativistic outflows.
@@ 5 Conclusions
The disk outflows are driven by a comparable contribution of vertical magnetic field lines threading the accretion disk (BP jets) and magnetic flux eruptions from regions of high magnetic pressure near the BH. Feedback from the cocoon, which is generated by the disk winds, induces a disk tilt that causes the outflows to wobble. With an isotropic-equivalent energy of $E_{\rm iso}\approx 10^{52}\,\,{\rm erg}$, these outflows explode the star, leaving merely $\sim 1\,\,{\rm M_{\odot}}$ for BH mass growth. After breakout from the star, the disk outflows will generate multiple types of emission across the spectrum: shock breakout, cooling envelope, ${}^{56}{\rm Ni}$ decay, and synchrotron emission. We find that their cooling emission will be detectable out to hundreds of Mpc by ULTRASAT and $D>{\rm Gpc}$ by UVEX and the Rubin Observatory. The detection rate of these events will provide valuable constraints on accretion disk physics and its formation rate in CCSNe.
The UV/optical cooling emission will be followed by ${}^{56}{\rm Ni}$ decay emission. Hydrodynamic simulations by Fujibayashi et al. (2023, 2024) suggest that the ${}^{56}{\rm Ni}$ emission is expected to last a few weeks, however, this result might change for magnetic disks [^2] Our hydrodynamic disks fail to launch such powerful outflows (see §3). This discrepancy may arise from the enforced axisymmetry in Fujibayashi et al. (2023, 2024), which requires verification through 3D calculations.. Additionally, recent neutrino–GRMHD simulations of collapsars indicate that strong magnetic fields in the accretion disk can facilitate the ejection of heavy $r$-process elements from the disk (Issa et al., 2024). This could potentially drive fast “kilonova” emission, enhancing the signal we find, and altering the ${}^{56}{\rm Ni}$ yields compared to those found in hydrodynamic simulations. Future work incorporating neutrino transport in GRMHD simulations is necessary to investigate the impact of disk cooling on our results and model all three optical signals (cooling, kilonova, and 56Ni decay) from first principles.